The Supply for product X is given by Qxs = -100 + 5Px

How much surplus do producers receive when Qx=253
Answer #1

Answer
the price is found as
Qx=253
Qx=-100+5Px
253=-100+5Px
5Px=353
Px=70.6
the inverse supply function is (to find out Y-intercept of the supply function)
Qx=-100+5Px
5Px=100+Qx
Px=20+0.2Qx

producer surplus =0.5*(price- Y-intercept of the supply function )*Qx
=0.5*(70.6-20)*252
=6375.6
the producer surplus is $6375.6